Car 1 has a full tank of gas, 13 gallons, but is leaking 2 gallons a week. Car 2 only has 8 gallons of gas but is also leaking at 1 gallon a week. When will the two cars have the same amount of gas in their tanks?
Solve the equation: X=?

Answer by ikleyn(52908)

.
13 - 2w = 8 - w  ====>

13 - 8 = 2w - w  ====>  5 = w  ====>  w = 5.


Answer. In 5 weeks.

In 5 weeks, both tanks will have 3 gallons of gas EACH.